please help me solve this problem,
Ten gallons of a 30% acid solution is obtained by mixing a 20% solution with a 50% solution. How much of each must be used?

Answer by scott8148(6628)

let x="gallons of 20%", so 10-x="gallons of 50%"

(20%)(x)+(50%)(10-x)=(30%)(10) Multiplying by 10 __ 2x+50-5x=30

subtracting 50 __ -3x=-20 __ dividing by -3 __ x=20/3

substituting __ 10-(20/3)=10/3